OpenStudy (warpedkitten):

ABCD ~ WXYZ. AD = 12, DC = 3, and WZ = 35. Find YZ. The figures are not drawn to scale.

OpenStudy (warpedkitten):

I did 12*35 and 3*x = 3x = 420 3x/3 = 420/3 = 140 but I also did 12*x and 3*35 105 = 12x 105/12 = 12x/12 =8.75 I'm not sure which is correct.

OpenStudy (warpedkitten):